OpenStudy (he66666):

Linear algebra: subspaces? Is the given subset of R^3 a subspace? The set of all vectors of the form [a; b; c], where a+2b-c=0 is a subspace. Answer: Yes I don't get why it is a subspace because it satisfies the closure properties? Can someone please help me?

OpenStudy (he66666):

it doesn't satisfy**

OpenStudy (turingtest):

\[\vec v=a+2b-c=0\]\[\vec u=x+2y-z=0\]add them together\[\vec v+\vec u=a+2b-c+(x+2y-z)=0\]\[\implies (a+x)+2(b+y)-(c+z)=0\]

OpenStudy (turingtest):

\[\vec w=(a+x)+2(b+y)-(c+z)=0\]let\[\begin{array}aa+x=p\\b+y=q\\c+z=r\end{array}\]then\[\vec w=p+2q-r=0\in W\]therefor closed under addition

OpenStudy (turingtest):

are you ok with that? the scalar part is trivial

OpenStudy (he66666):

Yes, I understand it now. Thanks!

OpenStudy (turingtest):

welcome :)